An electric dipole of moment P is lying along a uniform electric field E. The work done in rotating the dipole by 90° is




Grade:10

2 Answers

Amit Saxena
10 years ago

(c)

Work done in rotating a dipole , =pE(1-cosq)

If q = 90°, work done = pE (1- 0) = pE

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Workdone = ΔpE
pE = −p​E
pEf​ = 0
pEi​ = −pE
ΔpE = pE

Thanks and Regards